OpenStudy (anonymous):

find the area of a parallelogram that has a base of 2 1/2 ft and a height of 1 1/4 ft

OpenStudy (anonymous):

25/8

OpenStudy (anonymous):

yea 25/8

OpenStudy (anonymous):

area of ||gm = base x ht = 5/2 x 5/4 = 25/8

OpenStudy (anonymous):

now every single user write the same answer again..its a trend

OpenStudy (anonymous):

so is it 3 1/8

OpenStudy (anonymous):

yes